A second number is 6 less than a first number. A third number is twice the first number. If the sum of the three numbers is 306, find the numbers.
Found 2 solutions by solver91311, oberobic:
Answer by solver91311(24713)



Let represent the first number, then the second number must be and the third number must be . The sum is 306:



Solve for , then calculate , and

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by oberobic(2304)

x = first number
.
y = second number is 6 less than x
y = x-6
.
z = third number is 2 x
z = 2*x
.
x + y + z = 306
.
Substitute
.
x + (x-6) + 2x = 306
4x = 312
x = 78
.
y = 78 -6
y = 72
.
z = 2x
z = 2*78
z = 156
.
Always check your work.
78+72+156 = ??
= 306
Correct.
.
Answer: The three numbers are 78, 72, and 156.
.
Done.